. ? Express the distance, d, from Co to a point (x,y) ony = V/ as a function of x. (10)
Answer #1

५-2 (a) 4 /-y-+(1-0 2 ( py distayi ___ = 16-1)-+1- (My =12) ज-+ (1)2 (110) = 22-22+1+2c - --+ id=5a²_ath which is a function